Calculate the pressure of 10 mole NH3 gas in a10 L vessel at 0oC (273K) for both condition a) when the gas is ideal b) when the gas is real
R = 0.0821 L.atm.mol K a = 4.2L2.atm/mol2 b = 0.037 L/mol
1
Expert's answer
2015-11-02T12:33:11-0500
a) The ideal gas pressure can be calculated using the ideal gas law:
P = nRT/V
P = 10x8.314x273/0.01 = 2269848 Pa

b) The real gas pressure can be calculated using the Van der Waals law:
(p+n2a/V2) (V - nb) = nRT